OpenStudy (anonymous):

The formula to determine energy is E=1/2mv^2. What is the formula solved for v? options are 1. v= +sq. root of m/2E 2. v= 2E- m 3. v= + sq. root 2E-m 4. v= + sq. root 2E/m

OpenStudy (anonymous):

4 is the correct5 ans

OpenStudy (anonymous):

e=1/2mv^2 2E=mv^2 2E/m=v^2 root under 2E/m=v

OpenStudy (anonymous):

thank you (: